xx TestDerivs

TestDerivs

Liwei Ji <ljsma@rit.edu>

July 27 2023

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11 derivxorder 10 Scope: private01  INT

11 Description: Order of spatial finite differencing

11 Range 10              01  Default: 4
11 2  10 Second order finite difference
11 4  10 Fourth order finite difference
11 6  10 Sixth order finite difference

11

11 kxx  10 Scope: private01  REAL

11 Description: par for polynomial

11 Range 10              01  Default: 0.0

11 *:*  10

11

11 kxy  10 Scope: private01  REAL

11 Description: par for polynomial

11 Range 10              01  Default: 0.0

11 *:*  10

11

11 kyz  10 Scope: private01  REAL

11 Description: par for polynomial

11 Range 10              01  Default: 0.0

11 *:*  10

11

11 refinedxradius10 Scope: private01  REAL

11 Description: size of the refined region at the center

11 Range 10              01  Default: 0.25
11 (0:*  10 any positive size

11

7 Interfaces

General Implements: testderivs

Inherits: carpetxregrid Grid Variables

7.0.1 PRIVATE GROUPS

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  chi  10  chi  01  centering 11 centering={0 0 0}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  dchi  10                    01  centering 11 centering={0 0 0}
11                    10  dxchi  01  compact 11 0
11                    10  dychi  01 description 11 1st derivs of test grid function
11                    10  dzchi  01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  ddchi  10                    01  centering 11 centering={0 0 0}
11                    10  dxxchi  01  compact 11 0
11                    10  dxychi  01 description 11 2nd derivs of test grid function
11                    10  dxzchi  01 dimensions 11 3
11                    10  dyychi  01 distribution 11 DEFAULT
11                    10  dyzchi  01 group type 11 GF
11                    10  dzzchi  01  tags 11 checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11

Uses header: defs.hxx loopxdevice.hxx simd.hxx vect.hxx derivs.hxx

8 Schedule This section lists all the variables which are assigned storage by thorn CarpetX/TestDerivs. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11 Always: 11 x

11 chi  11 x

11 dchi  11 x
11 ddchi  11 x

11 x  11 x

11

Scheduled Functions

CCTKxPOSTINITIAL x testderivsxseterror x set up test grid x

11 x 11 Language: 11 c
11 x 11 Type:  11 function
11 x 11 Writes: 11 carpetxregrid::regridxerror(interior)

11

CCTKxINITIAL x testderivsxset x set up test data x

11 x 11 Language: 11 c
11 x 11 Sync:  11 chi
11 x 11 Type:  11 function
11 x 11 Writes: 11 chi(interior)

11

CCTKxPOSTREGRID x testderivsxsync x synchronize x

11 x 11 Language: 11 c
11 x 11 Options: 11 global
11 x 11 Sync:  11 chi
11 x 11 Type:  11 function

11

CCTKxPOSTSTEP x testderivsxcalcderivs x calculate derivs

11 x 11 Language: 11 c
11 x 11 Reads:  11 chi(everywhere)
11 x 11 Sync:  11 dchi
11 x 11 x  11 ddchi
11 x 11 Type:  11 function
11 x 11 Writes: 11 dchi(interior)
11 x 11 x  11 ddchi(interior)